A (g) + B(g) = C(g) + D(g)
Above equilibrium is established by taking A and B in a closed container. Initial concentration of A is twice the initial concentration of B. At equilibrium concentration of B and C are equal. Then find the equilibrium constant for the reaction, C(g) + D(g) = A(g) + B(g)

Grade:12

3 Answers

mohit yadav
11 years ago

the equilbrium constant for the reaction is 2 if the conc. of D is same as that of conc.of B.

Vikas TU
10 years ago

              A (g) + B(g)  =  C(g) + D(g)

t=0         2a        a            0         0

t=t       2a - x     a - x       x         x

 

At equilibrium concentration of B and C are equal.

a - x   =    x

a = 2x

x = a/2

Now,

at t=t 3a/2      a/2         a/2      a/2

 

K = [A] [B]/[C] [D]

K = [3a/2] [a/2]/ [a/2]2

K = 3a^2/4/a^2/4

K = 3

is te ansr.

 

Approve! plz

Fggg
4 years ago
Maderchod
Maderchod
Maderchod
Made
Maderchod
Maderchod
Maderchod
Maderchod
Maderchod
 
Maderchod
Maderchod
Maderchod